OpenStudy (newarf):

how could you rewrite the expreason (3 + 5)+8 using the associative property of addition

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (newarf):

no im confused

OpenStudy (skullpatrol):

Recall: when you add any three real numbers the grouping (or association) does not affect the result.

OpenStudy (skullpatrol):

( 3 + 5 ) + 8 means you are grouping the 3 and the 5 together first, right?

OpenStudy (newarf):

that's what I thought

OpenStudy (newarf):

I think its 5 + 3 + 8

OpenStudy (skullpatrol):

how would you group the 5 and the 8 together first?

OpenStudy (newarf):

is my answer right or wrong

OpenStudy (skullpatrol):

Nope. Keep the order of the numbers the same.

OpenStudy (skullpatrol):

3 + 5 + 8

OpenStudy (newarf):

8 + 3 + 5

OpenStudy (skullpatrol):

The property for changing the order of the numbers is called "commutative property" a + b = b + a

OpenStudy (newarf):

ok

OpenStudy (skullpatrol):

We want ( a + b ) + c = a + ( b + c )

OpenStudy (skullpatrol):

See how the a, b, and c stayed in order on both sides of the equation?

OpenStudy (newarf):

ok the answer I have that has that is 3 + (5 + 8)

OpenStudy (newarf):

is that right

OpenStudy (skullpatrol):

Yes :D

OpenStudy (newarf):

thank u so much

OpenStudy (skullpatrol):

Thanks for trying to learn!